# A Geometric View of Data Complexity: Efficient Local Intrinsic Dimension Estimation with Diffusion Models

### v1

@@ Appendix C Adapting FLIPD for DDPMs
DDPMs model the backward diffusion process (or denoising process) by modelling a network ${\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}\epsilon}:\mathbb{R}^{D}\times\{1,\ldots,{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}T}\}\to\mathbb{R}^{D}$ that takes in a noised-out point and outputs a residual that can be used to denoise. In particular, for every ${\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}t}$:
${\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}x_{t}}+{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}\epsilon}({\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}x_{t}},{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}t})\sim{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}p}({\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}x_{t-1}}|{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}x_{t}}).$ (75)
Song et al. [56] show that one can draw an equivalence between the network ${\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}\epsilon}(\cdot,{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}t})$ and the score network (see their Appendix B). Here, we rephrase the connections in a more explicit manner where we note that:
@@ D.4 In-depth Analysis of the Synthetic Benchmark
Among model-free methods, LPCA and ESS show good performance in low dimensions, with LPCA being exceptionally good in scenarios where the manifold is affine. As we see in Table 6, while model-free methods produce reliable estimates when $D$ is small, as $D$ increases the estimates become more unreliable. In addition, we include average LID estimates in Table 8 and see that all model-free baselines underestimate intrinsic dimensionality to some degree, with LPCA and MLE being particularly drastic when $D>100$. We note that ESS performs relatively well, even beating model-based methods in some $800$-dimensional scenarios. However, we note that both the $\mathcal{C}$ values in Table 7 and MAE values in Tables 6 and 8 suggest that none of these estimators effectively estimate LID in a pointwise manner and cannot rank data by LID as effectively as FLIPD.
@@ D.6 Improving the NB Estimators with kneedle
We recall that the NB estimator requires computing $\operatorname{rank}S(x)$, where $S(x)$ is a $K\times D$ matrix formed by stacking the scores $\hat{s}(\cdot,t_{0})$. We set $t_{0}=0.01$ as it provides the most reasonable estimates. To compute $\operatorname{rank}S(x)$ numerically, Stanczuk et al. [57] perform a singular value decomposition on $S(x)$ and use a cutoff threshold $\tau$ below which singular values are considered zero. Finding the best $\tau$ is challenging, so Stanczuk et al. [57] propose finding the two consecutive singular values with the maximum gap. Furthermore, we see that sometimes the top few singular values are disproportionately higher than the rest, resulting in severe overestimations of the LID. Thus, we introduce an alternative algorithm to determine the optimal $\tau$. For each $\tau$, we estimate LID by thresholding the singular values. Sweeping 100 different $\tau$ values from 0 to $1000$ at a geometric scale (to further emphasize smaller thresholds) produces estimates ranging from $D$ (keeping all singular values) to 0 (ignoring all). As $\tau$ varies, we see that the estimates plateau over a certain range of $\tau$. We use kneedle to detect this plateau because the starting point of a plateau is indeed a knee in the curve. This significantly improves the baseline, especially in high dimensions: see the third column of Tables 6, 8, and 7 compared to the second column.
@@ Appendix E Image Experiments
E.1 FLIPD Curves and DM Samples
Figures 9 and 10 show DM-generated samples and the associated LID curves for $4096$ samples from the datasets. In Figure 10, the FLIPD curves with MLPs, which have clearly discernible knees as predicted by the theory, are more reasonable than those generated with UNets. The fact that FLIPD estimates are worse for UNets is surprising given that MLPs produce worse-looking samples, as shown in Figure 9. However, comparing the first and second rows of Figure 9, it also becomes clear that MLP-generated images still match some characteristics of the true dataset, suggesting that they may still be capturing the image manifold in useful ways. To explain the surprisingly poor FLIPD estimates of UNets, we hypothesize that the convolutional layers in the UNet provide some inductive biases which, while helpful to produce visually pleasing images, might also encourage the network to over-fixate on high-frequency features which are not visually perceptible. For example, Kirichenko et al. [33] showed how normalizing flows over-fixate on these features, and how this can cause them to assign large likelihoods to out-of-distribution data, even when the model produces visually convincing samples. We hypothesize that a similar underlying phenomenon might be at play here, and that DMs with UNets might be over-emphasizing “high-frequency” directions of variation in their LID estimates, even if they learn the semantic ones and thus produce pleasing images. However, an exploration of this hypothesis is outside the scope of our work, and we highlight once again that FLIPD remains a useful measure of complexity when using UNets.

### v2

@@ Appendix C Adapting FLIPD for DDPMs
DDPMs model the backward diffusion process (or denoising process) by modelling a network ${\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}\epsilon}:\mathbb{R}^{D}\times\{1,\ldots,{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}T}\}\to\mathbb{R}^{D}$ that takes in a noised-out point and outputs a residual that can be used to denoise. Song et al. [57] show that one can draw an equivalence between the network ${\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}\epsilon}(\cdot,{\color[rgb]{0.64705882,0.43921569,0.2}\definecolor[named]{pgfstrokecolor}{rgb}{0.64705882,0.43921569,0.2}t})$ and the score network (see their Appendix B). Here, we rephrase the connections in a more explicit manner where we note that:
@@ D.4 In-depth Analysis of the Synthetic Benchmark
Among model-free methods, LPCA and ESS show good performance in low dimensions, with LPCA being exceptionally good in scenarios where the manifold is affine. As we see in Table 7, while model-free methods produce reliable estimates when $D$ is small, as $D$ increases the estimates become more unreliable. In addition, we include average LID estimates in Table 9 and see that all model-free baselines underestimate intrinsic dimensionality to some degree, with LPCA and MLE being particularly drastic when $D>100$. We note that ESS performs relatively well, even beating model-based methods in some $800$-dimensional scenarios. However, we note that both the $\mathcal{C}$ values in Table 8 and MAE values in Tables 7 and 9 suggest that none of these estimators effectively estimate LID in a pointwise manner and cannot rank data by LID as effectively as FLIPD.
@@ D.6 Improving the NB Estimators with kneedle
We recall that the NB estimator requires computing $\operatorname{rank}S(x)$, where $S(x)$ is a $K\times D$ matrix formed by stacking the scores $\hat{s}(\cdot,t_{0})$. We set $t_{0}=0.01$ as it provides the most reasonable estimates. To compute $\operatorname{rank}S(x)$ numerically, Stanczuk et al. [58] perform a singular value decomposition on $S(x)$ and use a cutoff threshold $\tau$ below which singular values are considered zero. Finding the best $\tau$ is challenging, so Stanczuk et al. [58] propose finding the two consecutive singular values with the maximum gap. Furthermore, we see that sometimes the top few singular values are disproportionately higher than the rest, resulting in severe overestimations of the LID. Thus, we introduce an alternative algorithm to determine the optimal $\tau$. For each $\tau$, we estimate LID by thresholding the singular values. Sweeping 100 different $\tau$ values from 0 to $1000$ at a geometric scale (to further emphasize smaller thresholds) produces estimates ranging from $D$ (keeping all singular values) to 0 (ignoring all). As $\tau$ varies, we see that the estimates plateau over a certain range of $\tau$. We use kneedle to detect this plateau because the starting point of a plateau is indeed a knee in the curve. This significantly improves the baseline, especially in high dimensions: see the third column of Tables 7, 9, and 8 compared to the second column.
@@ Appendix E Image Experiments
E.1 FLIPD Curves and DM Samples: A Surprising Result
Figures 9 and 10 show DM-generated samples and the associated LID curves for $4096$ samples from the datasets. In Figure 10, the FLIPD curves with MLPs have clearly discernible knees at which the estimated LID averages around a positive number, as expected. When changing the architecture of the DMs from MLPs to UNets, however, the curves either lack knees or obtain them at negative values of LID. This makes it harder to set $t_{0}$ automatically with kneedle when using UNets, and thus makes the estimator non-robust to the choice of architecture. Note that our theory ensures that $\text{FLIPD}(x,t_{0})$ converges to $\text{LID}(x)$ as $t_{0}\rightarrow 0$ when $\hat{s}$ is an accurate estimator of the true score function; indeed, it does not make any claims about the behaviour of FLIPD when $t_{0}$ is large or when the model fit is poor. Yet, the fact that the FLIPD estimates obtained from UNets are less reliable than the ones from MLPs is somewhat surprising, especially given that DMs with MLP backbones clearly produce worse-looking samples (shown in Figure 9), suggesting that while the model fit is better in the former case, the LID estimates are worse. Throughout the paper, we have argued that, notwithstanding this, FLIPD estimates derived from both architectures offer useful measurements of relative complexity, and can still effectively rank data based on its complexity; refer to all the images in Figures 17, 18, 19, 20, 21, 22, 23, and 24. However, this discrepancy between FLIPD estimates – in absolute terms – when altering the model architecture is counterintuitive and requires further treatment. In what follows, we propose hypotheses for the significant variation in FLIPD estimates when transitioning from MLP to UNet architectures, and we encourage further studies to investigate this issue and to adjust FLIPD to better suit DMs with state-of-the-art backbones.
While MLP architectures do not produce visually pleasing samples, comparing the first and second rows of Figure 9 side-by-side, it also becomes clear that MLP-generated images still match some characteristics of the true dataset, suggesting that they may still be capturing the image manifold in useful ways. To explain the unstable FLIPD estimates of UNets, we hypothesize that the convolutional layers in the UNet provide some inductive biases which, while helpful to produce visually pleasing images, might also encourage the network to over-fixate on high-frequency features. More specifically, score functions can be interpreted as providing a denoiser which removes noise added during the diffusion process; our hypothesis is that UNets are particularly good at removing noise along high-frequency directions but not necessarily along low-frequency ones. Inductive biases or modelling errors which make the learned score function $\hat{s}$ deviate from $s$ might significantly alter the behaviour of FLIPD estimates, even if they result in visually pleasing images; for a more formal treatment of the effect of this deviation, please refer to the work of Tempczyk et al. [62], who link modelling errors to LID estimation errors in LIDL.
The aforementioned inductive biases have been found to produce unintuitive behaviours in other settings as well. For example, Kirichenko et al. [34] showed how normalizing flows over-fixate on these high-frequency features, and how this can cause them to assign large likelihoods to out-of-distribution data, even when the model produces visually convincing samples. We hypothesize that a similar underlying phenomenon might be at play here, and that DMs with UNets might be over-emphasizing high-frequency directions of variation in their LID estimates. On the other hand, MLPs do not inherently distinguish between the high- and low-frequency directions and can therefore more reliably estimate the manifold dimension, even though their generated samples appear less appealing to the human eye. However, an exploration of this hypothesis is outside the scope of our work, and we highlight once again that FLIPD remains a useful measure of relative complexity when using either UNets or MLPs.